OpenStudy (loser66):

We have \[\sum_{j =1}^n j = \dfrac{n(n+1)}{2}\] How about the limit from m to n? I meant \[\sum_{j =m}^n j=??\] Please, help

Parth (parthkohli):

\[= \sum_{1}^{n}j - \sum_{1}^{m-1}j\]

Parth (parthkohli):

sorry for lack of notation

imqwerty (imqwerty):

terms from m to n = n-m+1 so summation wuld be- \[\frac{ (n-m+1)(m+n) }{ 2 }\]

imqwerty (imqwerty):

number of terms =n-m+1 1st term =m and common diff=1 so just put this in the formula of summation of ap :)

OpenStudy (loser66):

Got it. thank you.

imqwerty (imqwerty):

np :)

OpenStudy (loser66):

In general, if we see the lower limit is not 1, when we see n, we replace by what? hehehe. I am silly, just want to generalize the formula like \(\sum_{j = 1}^n = \dfrac{n(n+1)}{2} \) the second n+ 1 is upper + lower limit, right? the first n is upper - lower +1 Hence for all lower limit, we just apply it, right?

OpenStudy (loser66):